What is/are the outlier/s in this data set?

 10, 32, 19, 8, 22, 29, 68,36,  14, 49, .


1
Expert's answer
2021-12-20T05:55:40-0500



the outlier is an extreme observation there are lower and upper outliers

lower outliner is an observation at the bottom and upper outlier is observed at the top.

mean=28.7

standard deviation = "\sqrt{ variance}"

variance ="\frac{\sum (x-\mu)^2}{n-1}"

=11640.69

standard deviation =35.964

25% =p(x<x)=0.25

=NORM.INV(0.25,28.7,35.964)

=4.443

75%=p(x<x)=0.75

=NORM.INV(0.75,,28.7,35.964)

= 52.9573

IQR=52.9573-4.443

=48.51

Lowe outlier =Q1-(1.5*48.51)=-68.3224

Upper outlier=Q3-(1.5*48.51)=125.727

from the data there are no outliers since we didn't have values from extreme ends.